# When Deepfakes Look Real: Detecting AI-Generated Faces with Unlabeled Data due to Annotation Challenges

### v2

@@ Abstract
Existing deepfake detection methods heavily depend on labeled training data. However, as AI-generated content becomes increasingly realistic, even human annotators struggle to distinguish between deepfakes and authentic images. This makes the labeling process both time-consuming and less reliable. Specifically, there is a growing demand for approaches that can effectively utilize large-scale unlabeled data from online social networks. Unlike typical unsupervised learning tasks, where categories are distinct, AI-generated faces closely mimic real image distributions and share strong similarities, causing performance drop in conventional strategies. In this paper, we introduce the Dual-Path Guidance Network (DPGNet), to tackle two key challenges: (1) bridging the domain gap between faces from different generation models, and (2) utilizing unlabeled image samples. The method features two core modules: text-guided cross-domain alignment, which uses learnable prompts to unify visual and textual embeddings into a domain-invariant feature space, and curriculum-driven pseudo label generation, which dynamically exploit more informative unlabeled samples. To prevent catastrophic forgetting, we also facilitate bridging between domains via cross-domain knowledge distillation. Extensive experiments on 11 popular datasets, show that DPGNet outperforms SoTA approaches by 6.3%, highlighting its effectiveness in leveraging unlabeled data to address the annotation challenges posed by the increasing realism of deepfakes. [^1] The code will be open-sourced upon publication.
Introduction
The rapid rise of deepfakes(Zhou et al. 2023; Yu et al. 2023; Zhang et al. 2023, 2024a; Pan et al. 2024; Zhang et al. 2024b; Liu, Ye, and Du 2024), especially in the form of face forgery, has emerged as a major challenge to media authenticity. Face forgery, which involves manipulating or generating human faces in images and videos, poses significant risks across various sectors, including politics, security, and entertainment. These generated media are often so realistic that they are nearly indistinguishable from real footage, eroding public trust in visual content and raising concerns about their potential misuse. As a result, detecting deepfakes, particularly face forgery, has become a critical area of research in artificial intelligence and digital forensics.
@@ Introduction
Traditional deepfake detection methods(Nguyen et al. 2024; Larue et al. 2023; Qiao et al. 2024; Lin et al. 2024a; Luo et al. 2023) rely heavily on labeled training data, where human annotators manually classify images or videos as either AI-generated or real. While these supervised methods have been effective to some extent, their limitations are becoming more apparent as deepfake generation techniques advance. As AI-generated faces become increasingly realistic, human annotators struggle to make accurate distinctions, rendering the labeling process both time-consuming and less reliable. This reliance on labeled data creates a significant bottleneck, limiting the scalability and practicality of current detection systems, especially as the volume of digital content continues to grow exponentially. Given these challenges, there is an urgent need for deepfake detection methods that can effectively leverage large-scale unlabeled data. Such approaches could alleviate the burden of manual annotation and enable more scalable systems. The vast amounts of unlabeled data available from online social networks offer a unique opportunity to train deepfake detection models without relying on expensive labeled datasets.
@@
Can traditional unsupervised learning methods handle this task? A key challenge in deepfake detection is that faces generated by different AI models closely mimic the distribution of real human faces and are often highly similar to each other. Unlike typical unsupervised learning tasks, where semantic categories are well-separated, faces produced by generative models share many common features with real faces, leading to significant overlap. This overlap complicates the task for traditional unsupervised methods, which rely on clearly defined categories to differentiate between real and fake. As a result, existing unsupervised learning(Bai et al. 2024; Yu, Huang, and Zhang 2025; Zhuang et al. 2022; Deng et al. 2025; Zhang et al. 2025) approaches struggle to capture the subtle differences between real and fake faces, resulting in lower performance and reduced effectiveness in practical applications.
@@
In this work, we introduce the Dual-Path Guidance Network(DPGNet), a novel framework designed to tackle the challenge introduced above. Unlike traditional methods that rely solely on labeled data, DPGNet combines two paths. First, we retain the original labeled data from traditional training settings as the source domain. The second path takes advantage of large-scale unlabeled data, often sourced from online social networks, which reflects the abundant real-world data available for training. DPGNet addresses two main challenges: (1) bridging the gap between labeled source data and the diverse, unlabeled data generated by different AI models, and (2) effectively utilizing large-scale unlabeled images. DPGNet consists of two key components: text-guided cross-domain alignment and curriculum-driven pseudo label generation. The first component uses learnable prompts to align visual and textual information into a shared, domain-independent feature space. This allows the model to better handle different types of deepfake faces while leveraging textual information. The second component mimics human learning by gradually incorporating and learning from more informative unlabeled samples. Through dynamic threshold supervision, it ensures the model focuses on the challenging samples.
@@
We conduct extensive experiments across 11 datasets, including both cross-domain and cross-method evaluations, to evaluate the effectiveness of DPGNet. Our results show that the proposed method outperforms SoTA methods, achieving a significant improvement of 6.3% in detection accuracy. These findings highlight the ability of our method to effectively leverage unlabeled data in real-world scenarios, overcoming the annotation challenges caused by the increasing realism of deepfakes and providing a scalable solution for face forgery detection in the age of AI-generated media. The main contributions are summarized as follows:
@@
Figure 2: Comparison of methods leveraging unlabeled data. CLIP-Source is trained on FF++ (Rossler et al. 2019a). CLIP+PL(x) variants are fine-tuned using UCDDP pseudo-labels with confidence thresholds $\geq 0.9$, $0.8$, and $0.7$. CLIP+DA uses domain alignment with UCDDP, a small subset sampled from the training sets of the five test datasets.
  • We are the first to effectively leverage unlabeled data for face forgery detection, addressing the growing annotation challenges posed by increasingly realistic AI-generated content and advanced generation techniques.
  • We introduce the DPGNet, a novel framework that combines two paths: one for bridging the gap between labeled source data and diverse, unlabeled data generated by different methods, and another for effectively utilizing large-scale unlabeled data to improve performance.
@@
  • Comprehensive experiments have validated the effectiveness of DPGNet across 11 popular datasets under various settings, demonstrating superior performance (+6.3%).
@@
Related Work
Our work focuses on leveraging unlabeled data in face forgery detection to bridge the domain gap between source and target domains, preserving prior knowledge while enhancing robustness to various forgery types. We review related works on generalizable deepfake image detection and domain adaptation, highlight their limitations and elaborate on our contributions.
Generalizable deep fake image detection. The emergence of new fake methods has become an important challenge that plagues detector research. To solve this problem, generalization has become the mainstream direction of current research. Some works design detectors by mining general fake clues that may exist in fake images. (Yan et al. 2023; Fu et al. 2025a; Huang et al. 2023; Dong et al. 2023) extract general fake features for learning by decoupling and reconstructing images or IDs. (Li et al. 2021; Luo et al. 2021; Liu et al. 2021a) trains detectors by mining the differences between fake and real images in the frequency domain. (Yan et al. 2024a) uses distillation learning between teacher and student encoders to enable student encoders to learn more general fake knowledge. Some works pursue generalization by simulating more fake methods through operations such as data augmentation/mixing. In recent studies, (Yan et al. 2024a; Yermakov, Cech, and Matas 2025; Cui et al. 2024) leverage the powerful feature extraction capabilities of large visual language models such as clip to achieve efficient generated image detection. (Cui et al. 2024) use adapters to guide clip to pay attention to fake clues. (Yan et al. 2024a) and (Yermakov, Cech, and Matas 2025) respectively proposed two different fine-tuning methods to unlock the potential of clips by exploring their weighted contextual relationships.
Unsupervised Domain Adaptation (UDA). Unsupervised domain adaptation aims to address the domain gap in detection by leveraging unlabeled target domain data. Early UDA methods (Long et al. 2015, 2017; Sun and Saenko 2016) focus on learning domain-invariant features by minimizing domain differences. For example, DAN (Long et al. 2015) uses the maximum mean difference to align domains, while CORAL (Sun and Saenko 2016) employs linear projection to match second-order statistics. Adversarial methods, such as DANN (Ganin and Lempitsky 2015) and CDAN (Long et al. 2018), use domain discriminators to align source and target feature distributions. However, these methods perform poorly in deep face forgery detection due to a key difference: unlike traditional classification tasks with clear semantic boundaries (e.g., ‘cat’ vs. ‘dog’), real and fake faces share the same high-level attributes (‘face’). The difference between them lies in the traces of artifacts (e.g., artifacts in the eye region or inconsistent skin texture) rather than different semantic categories. Pseudo-labeling strategies (Zhou et al. 2024)alleviate this by assigning labels to unlabeled data, but high-confidence pseudo-labels often prioritize simple samples with low value, leading to error propagation and ignoring challenging cases that are critical for robust detection. Some parameter efficient fine-tuning (PEFT) methods (Han et al. 2024), such as LoRA(Hu et al. 2022), achieve domain adaptation by fine-tuning the visual base model, but there is a risk of distorting the pre-trained knowledge (Cai et al. 2019; Tang, Chen, and Jia 2020), which may lead to low ranking (Yan et al. 2024b) in the feature space. Our method overcomes these limitations by integrating visual-language alignment and dynamic pseudo-labeling, effectively capturing various artifact patterns while preserving prior knowledge and ensuring robust generalization across domains.
Method
Problem Definition
@@
To emulate the diverse forgery techniques encountered in real-world scenarios, we constructed two composite datasets, UCDDP and UDF40, by sampling a small subset of images from the training sets of multiple cross-domain datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{K}\}$. This approach ensures diversity in unknown forgery methods and data distributions. Specifically, UCDDP encompasses images generated by various unknown forgery techniques, exhibiting significant distributional variations across samples. In contrast, UDF40 includes forgery methods distinct from those in the source domain while maintaining a consistent data distribution with the source domain. The task is formalized as minimizing the expected classification error on the target domain:
@@
Inspired by this, we propose using text clues as a bridge to coordinate source-target domain knowledge, improve distribution shift, and introduce a curriculum learning strategy to dynamically integrate high-value difficult samples. The DPGNet consists of two stages: source domain pre-training and joint domain adaptation. In the first stage, to establish text-guided domain alignment, given the pre-processed source domain image $x_{s}\in\mathcal{D}_{s}$, we use the visual encoder $E_{v}$ to extract semantically rich features $z_{s}^{f}\in\mathbb{R}^{256\times 1024}$ from it, and align $z_{s}^{f}$ with the real/fake specific text vectors $e_{t}^{f}\in\mathbb{R}^{768}$ generated by the hint learning module. This alignment is optimized through a composite constraint $\mathcal{L}_{\text{Source}}$, ensuring that $z_{s}^{f}$ captures robust, category-independent semantic representations, while $f_{\text{real}},f_{\text{fake}}$ encodes domain-invariant features of real and fake samples.
In the second phase, for unlabeled images from the target domain $x_{t}\in\mathcal{D}_{t}$, the DPGNet extracts features using a fine-tuned encoder $E_{v}$ to obtain $z_{t}^{f}$, perform classification inference, and pad them according to a high confidence threshold $\lambda_{tf}$, thus obtaining a feature base for the target domain $B_{\text{real}},B_{\text{fake}}$. To generate reliable pseudo labels, we measure the feature distances between $z_{t}^{f}$ and $B_{\text{real}},B_{\text{fake}}$ in the feature library and combine them with the classifier predictions to obtain pseudo labels. To improve the quality of pseudo labels, we introduce a curriculum learning strategy to dynamically adjust the screening threshold $\lambda_{lt}$ to merge challenging high-value samples. To mitigate catastrophic forgetting, we apply cross-domain augmentation to source domain features $z_{s}^{f}$ and employ knowledge distillation to align representations of $\mathcal{D}_{s}$ and $\mathcal{D}_{u}$, enabling joint training and enhancing robust generalization across domains.
Text-Guided Cross-Domain Alignment
In this module, the visual encoder is jointly trained with learnable text prompts. Drawing inspiration from prior work (Yermakov, Cech, and Matas 2025), we selectively fine-tune the layer normalization parameters of the first 24 Transformer layers to preserve pre-trained knowledge. For a source domain image $x_{s}\in\mathcal{D}_{s}$, the visual encoder $E_{v}$ extracts visual embedding features $z_{s}^{f}\in\mathbb{R}^{256\times 1024}$, which are processed by a classification head $h(\cdot)$ to produce the prediction $\hat{y}_{s}=h(z_{s}^{f})$. This process enables $E_{v}$ to effectively capture forgery-related features, which can be formulated as:
@@
Learnable Text Prompts. We introduce two trainable text prompts, initialized as ‘real face photo’ and ‘deep fake face photo’, parameterized as $v_{\text{real}}\in\mathbb{R}^{d}$ and $v_{\text{fake}}\in\mathbb{R}^{d}$, where $d$ is the text embedding dimension. These prompts are fed into the CLIP text encoder $E_{t}$ to generate concept embeddings:
$f_{\text{real}}^{f}=f_{t}(v_{\text{real}}),\quad f_{\text{fake}}^{f}=f_{t}(v_{\text{fake}})$ (3)
where these embeddings serve as semantic anchors for real and fake categories, capturing domain-invariant, independent concepts that transcend source-target distribution shifts.
Visual-Text Alignment. To ensure that the semantics encoded by the visual feature $z_{i}^{f}$ are compatible with $e_{y_{i}}^{f}$, we perform visual-text alignment:
$\mathcal{L}_{\text{alig}}=-\log\frac{\exp(\text{sim}(z_{i}^{f},f_{y_{i}})/\tau)}{\exp(\text{sim}(z_{i}^{f},f_{\text{real}})/\tau)+\exp(\text{sim}(z_{i}^{f},f_{\text{fake}})/\tau)}$ (4)
@@
$\mathcal{L}_{\text{con}}=\mathbb{E}_{x_{i}\sim\mathcal{D}_{i}}\left[-\text{sim}(z_{i}^{f},f_{y_{i}})+\text{sim}(z_{i}^{f},f_{y_{i}})\right]$ (5)
where $f_{y_{i}}$ is the text embedding of the opposite class. This process prioritizes task-relevant semantic features $\mathcal{S}_{c}\subseteq\mathcal{F}_{v}$ (related to authenticity) over domain-specific features $\mathcal{A}_{d}$.
Curriculum-Driven Pseudo Label Generation

### v3

# Leveraging Unlabeled Data from Unknown Sources via Dual-Path Guidance for Deepfake Face Detection

@@ Abstract
Existing deepfake detection methods heavily rely on static labeled datasets. However, with the proliferation of generative models, real-world scenarios are flooded with massive amounts of unlabeled fake face data from unknown sources. This presents a critical dilemma: detectors relying solely on existing data face generalization failure, while manual labeling for this new stream is infeasible due to the high realism of fakes. A more fundamental challenge is that, unlike typical unsupervised learning tasks where categories are clearly defined, real and fake faces share the same semantics, which leads to a decline in the performance of traditional unsupervised strategies. Therefore, there is an urgent need for a new paradigm designed specifically for this scenario to effectively utilize these unlabeled data. Accordingly, this paper proposes a dual-path guided network (DPGNet) to address two key challenges: (1) bridging the domain differences between faces generated by different generative models; and (2) utilizing unlabeled image samples. The method comprises two core modules: text-guided cross-domain alignment, which uses learnable cues to unify visual and textual embeddings into a domain-invariant feature space; and curriculum-driven pseudo-label generation, which dynamically utilizes unlabeled samples. Extensive experiments on multiple mainstream datasets show that DPGNet significantly outperforms existing techniques,, highlighting its effectiveness in addressing the challenges posed by the deepfakes using unlabeled data.  [^1] The code is in the supplementary material and will be open-sourced after publication.
1 Introduction
The proliferation and accessibility of generative models [zhou2023generative, yu2023reinforcement, zhang2023encrypted, zhang2024dibad, pan2024few, zhang2024few, liu2024learning] has led to an unprecedented surge in deepfakes, particularly in the form of face forgery. As a response, the dominant research paradigm [huang2025sida, liu2024forgerygpt, xu2024fakeshield, wang2025mllm, yang2025heie, peng2025mllm, guo2025rethinking] has relied on supervised learning. These methods are trained on static, curated datasets which are constructed using a limited, known set of forgery techniques. While effective in-domain, this reliance on known generators creates a fundamental vulnerability and is increasingly misaligned with the current media ecosystem.
@@ 1 Introduction
To address this vulnerability, much existing research has focused on improving generalization [laa-net, qiao2024fully, lin2024preserving, cfm, effort, ForensicsAdapter], aiming to create detectors that are robust to unseen forgery types. However, we posit that this defense-only approach overlooks a crucial opportunity. The reality of online platforms is a massive, continuous stream of unlabeled content where the generative source is unknown. We believe this data stream, rather than being merely an obstacle, can be actively leveraged to enhance detection. This perspective, however, leads to a critical dilemma: the unknown source reality means the data is unlabeled. The seemingly straightforward alternative of manually labeling this unlabeled data is fundamentally impractical. The high realism of modern fakes makes any such attempt both time-consuming and unreliable. This highlights an urgent need for a new paradigm: one that can effectively leverage this unlabeled and unknown-source data stream.
@@
Can traditional unsupervised learning methods handle this task? A key challenge in deepfake detection is that faces generated by different AI models closely mimic the distribution of real human faces and are often highly similar to each other. Unlike typical unsupervised learning tasks, where semantic categories are well-separated, faces produced by generative models share many common features with real faces, leading to significant overlap. This overlap complicates the task for traditional unsupervised methods, which rely on clearly defined categories to differentiate between real and fake. As a result, existing unsupervised learning [bai2024prompt, yu2025feature, zhuang2022uia, deng2025multi, zhang2025link, unsurpervised] approaches struggle to capture the subtle differences between real and fake faces, resulting in lower performance and reduced effectiveness in practical applications.
@@
In this work, we introduce the Dual-Path Guidance Network(DPGNet), a novel framework designed to tackle the challenge introduced above. Unlike traditional methods that rely solely on labeled data, DPGNet combines two paths. First, we retain the original labeled data from traditional training settings as the source domain. The second path takes advantage of large-scale unlabeled data, often sourced from online social networks, which reflects the abundant real-world data available for training. DPGNet addresses two main challenges: (1) bridging the gap between labeled source data and the diverse, unlabeled data generated by different AI models, and (2) effectively utilizing large-scale unlabeled images. DPGNet consists of two key components: text-guided cross-domain alignment and curriculum-driven pseudo label generation. The first component uses learnable prompts to align visual and textual information into a shared, domain-independent feature space. This allows the model to better handle different types of deepfake faces while leveraging textual information. The second component mimics human learning by gradually incorporating and learning from more informative unlabeled samples. Through dynamic threshold supervision, it ensures the model focuses on the challenging samples. To prevent the loss of previously learned information, we also introduce the cross-domain feature enhancement. This ensures that the source domain’s representation is bridged while adapting to new data.
@@
We conduct extensive experiments across multiple datasets, including both cross-domain and cross-method evaluations, to evaluate the effectiveness of DPGNet. Our results show that the proposed method outperforms SoTA methods, achieving a significant improvement of detection in detection accuracy. These findings highlight the ability of our method to effectively leverage unlabeled data in real-world scenarios, overcoming the annotation challenges caused by the increasing realism of deepfakes and providing a scalable solution for face forgery detection. The main contributions are summarized as follows:
@@
Figure 2: Comparison of methods leveraging unlabeled data. CLIP-Source is trained on FF++ [ff++]. CLIP+PL(x) variants are fine-tuned using UCDDP pseudo-labels with confidence thresholds $\geq 0.9$, $0.8$, and $0.7$. CLIP+DA uses domain alignment with UCDDP  1. See Section  3.1 for the details of the dataset.
  • We are the first to propose a paradigm for effectively utilizing unlabeled data for face forgery detection, shifting from a generalization used only for defense to actively utilizing unlabeled data streams from unknown sources.
  • We propose DPGNet, a novel dual-path framework that synergizes text-guided domain alignment and curriculum-driven pseudo-labeling to effectively learn from both labeled and unlabeled data streams.
@@
  • Comprehensive experiments have validated the effectiveness of DPGNet in various settings on multiple popular datasets and demonstrated its superior performance.
2 Related Work
Our work focuses on leveraging unlabeled data from unknown generators in face forgery detection to bridge the domain gap between source and target domains. We review related works on generalizable deepfake detection and traditional domain adaptation, highlight their limitations, and elaborate on our contributions.
Generalizable deep fake image detection. To address the emergence of new forgery methods, generalization has become a mainstream direction. Early works focused on mining general artifacts, such as frequency-domain discrepancies [li2021frequency, luo2021generalizing, liu2021spatial], or used reconstruction, decoupling, and distillation to learn shared fake features [yan2023ucf, fu2025exploring, huang2023implicit, dong2023implicit, lsda]. In recent studies, the field has shifted to leveraging the powerful features of V-LMs like CLIP, using adapters [ForensicsAdapter] or specialized fine-tuning [effort, 2025unlocking, guo2025rethinking]. Concurrently, MLLMs (Multimodal Large Language Models) are also being actively explored, with a growing number of methods [huang2025sida, xu2024fakeshield, wang2025mllm, yang2025heie, guo2025rethinking] leveraging the strong zero-shot reasoning and generalization capabilities inherited from their large-scale pre-training to assess image.
Unsupervised Domain Adaptation (UDA). UDA aims to address the domain gap using unlabeled target data. Methods in this area range from traditional adversarial alignment [long2017conditional, ganin2015unsupervised] to more recent techniques in knowledge transfer [ma2025steady] and adaptive alignment [he2025differential]. However, these methods perform poorly in deep face forgery detection, as the task fundamentally violates the core assumptions of UDA in two ways. First, the very definition of a domain is misaligned. In traditional UDA, domains have clear stylistic differences (e.g., photos vs. paintings). In forgery detection, however, different domains all appear as faces, rendering the inter-domain differences subtle and non-semantic. Second, the categories themselves are semantically inseparable. Unlike traditional classification tasks (e.g., cat vs. dog), real and fake faces share the identical high-level attribute of face. The distinction lies in subtle artifact traces, not in different semantic classes.
This dual challenge of non-semantic domains and semantically-identical categories invalidates the core premise of traditional feature alignment methods. To bypass this alignment failure, pseudo-labeling strategies [unsurpervised] are often employed. However, high-confidence pseudo-labels often prioritize simple samples with low value, leading to error propagation and ignoring challenging cases that are critical for robust detection. Some parameter efficient fine-tuning (PEFT) methods [PEFT], such as LoRA [lora], achieve domain adaptation by fine-tuning the visual base model, but there is a risk of distorting the pre-trained knowledge [cai2019learning, tang2020unsupervised], which may lead to low ranking [effort] in the feature space. Our method overcomes these limitations by integrating visual-language alignment and dynamic pseudo-labeling, effectively capturing various artifact patterns while preserving prior knowledge and ensuring robust generalization across domains.
@@ 2 Related Work
3 Methodology
3.1 Problem Definition
@@
To emulate the diverse forgery techniques encountered in real-world scenarios, we constructed two composite datasets, UCDDP and UDF40, by sampling a small subset of images from the training sets of multiple cross-domain datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{K}\}$. This approach ensures diversity in unknown forgery methods and data distributions. Specifically, UCDDP encompasses images generated by various unknown forgery techniques, exhibiting significant distributional variations across samples. In contrast, UDF40 includes forgery methods distinct from those in the source domain while maintaining a consistent data distribution with the source domain. The task is formalized as minimizing the classification error on the target domain:
@@ 3.2 Framework Overview
Inspired by this, we propose using text clues as a bridge to coordinate source-target domain knowledge, improve distribution shift, and introduce a curriculum learning strategy to dynamically integrate high-value difficult samples. The DPGNet consists of two stages: source domain pre-training and joint domain adaptation. In the first stage, to establish text-guided domain alignment, given the pre-processed source domain image $x_{s}\in\mathcal{D}_{s}$, we use the visual encoder $E_{v}$ to extract semantically rich features $z_{s}^{f}\in\mathbb{R}^{256\times 1024}$ from it, and align $z_{s}^{f}$ with the real/fake specific text vectors $e_{t}^{f}\in\mathbb{R}^{768}$ generated by the hint learning module. This alignment is optimized through a composite constraint $\mathcal{L}_{\text{Source}}$, ensuring that $z_{s}^{f}$ captures category-independent semantic representations, while $f_{r},f_{f}$ encodes domain-invariant features of real and fake samples.
In the second phase, for unlabeled images from the target domain $x_{t}\in\mathcal{D}_{t}$, the DPGNet extracts features using a fine-tuned encoder $E_{v}$ to obtain $z_{t}^{f}$, perform classification inference, and pad them according to a high confidence threshold $\lambda_{tf}$, thus obtaining a feature base for the target domain $B_{\text{real}},B_{\text{fake}}$. To generate reliable pseudo labels, we measure the feature distances between $z_{t}^{f}$ and $B_{\text{real}},B_{\text{fake}}$ in the feature library and combine them with the classifier predictions to obtain pseudo labels. To improve the quality of pseudo labels, we introduce a curriculum learning strategy to dynamically adjust the screening threshold $\lambda_{lt}$ to merge challenging high-value samples. To mitigate catastrophic forgetting, we apply cross-domain augmentation to source domain features $z_{s}^{f}$ and employ knowledge distillation to align representations of $\mathcal{D}_{s}$ and $\mathcal{D}_{u}$, enabling joint training and enhancing generalization across domains.
3.3 Text-Guided Cross-Domain Alignment
In this module, the visual encoder is jointly trained with learnable text prompts. Drawing inspiration from prior work [2025unlocking], we selectively fine-tune the layer normalization parameters of the first 24 Transformer layers to preserve pre-trained knowledge. For a source domain image $x_{s}\in\mathcal{D}_{s}$, the visual encoder $E_{v}$ extracts visual embedding features $z_{s}^{f}\in\mathbb{R}^{256\times 1024}$, which are processed by a classification head $h(\cdot)$ to produce the prediction $\hat{y}_{s}=h(z_{s}^{f})$. This process enables $E_{v}$ to effectively capture forgery-related features, which can be formulated as:
@@ 3.3 Text-Guided Cross-Domain Alignment
Learnable Text Prompts. We introduce two trainable text prompts, initialized as ”real face photo” and ”deep fake face photo”, parameterized as $v_{r}\in\mathbb{R}^{d}$ and $v_{f}\in\mathbb{R}^{d}$, where $d$ is the text embedding dimension. These prompts are fed into the CLIP text encoder $E_{t}$ to generate concept embeddings:
$f_{r}=f_{t}(v_{r}),\quad f_{f}=f_{t}(v_{f})$ (3)
where these embeddings serve as semantic anchors for real and fake categories, capturing domain-invariant concepts that transcend source-target distribution shifts.
Visual-Text Alignment. To ensure that the semantics encoded by the visual feature $z_{i}^{f}$ are compatible with the text embedding $f_{y_{i}}$, we perform visual-text alignment:
$\mathcal{L}_{\text{alig}}=-\log\frac{\exp(\text{sim}(z_{i}^{f},f_{y_{i}})/\tau)}{\sum_{c\in\{r,f\}}\exp(\text{sim}(z_{i}^{f},f_{c})/\tau)}$ (4)
@@
$\mathcal{L}_{{con}}=\mathbb{E}_{x_{i}\sim\mathcal{D}_{i}}\left[-\text{sim}(z_{i}^{f},f_{y_{i}})+\text{sim}(z_{i}^{f},\textbf{$f_{\neg y_{i}}$})\right]$ (5)
where $f_{\neg y_{i}}$ is the text embedding of the opposite class. This process prioritizes task-relevant semantic features $\mathcal{S}_{c}\subseteq\mathcal{F}_{v}$ (related to authenticity) over domain-specific features $\mathcal{A}_{d}$.
3.4 Curriculum-Driven Pseudo Label Generation